day shift Jan 30,2000


Subject: day shift Jan 30,2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:45:18 CST


Shift Summary for day shift, 01/30/00

On Shift: Mike Arenton, Albert Erwin

Hours of beam delivered at all intensities:4.5
Hours of beam delivered at nominal intensity: 4.5
Hours of KAMI data 4.0, beam still being tuned

Accounting of lost time:tev quenches,

===============================================================

Runs ended this shift:

Run 15685 long muon run (about 9 hours)
many short runs for various tests were done
===============================================================

Tapes brought to FCC this shift: none

Accesses during shift: none

===============================================================

Description of Shift:

run spill time comment
----------------------------------------------------
            06:50 tev quench

            07:26 beam returns
            09:25 quench
            10:08 beam returns
                   Rick Ford begins beam tuning.
            11:15 end long muon run 15685
            11:50 quench
 15686 12:00 KAMI_HADRON run, notice that the mysterious
                   low energy peak in UCAL is gone, realize that
                   now we have the sweepers on.
 15688 12:20 short run to get cosmics to tape for
                   background subtraction.
            14:10 beam returns, claim is that quench problem
                   is fixed.
 15689 one spill with sweepers off, low E peak reappears.
            14:20 turn on NM4AN, had problems turning it on, OPS fix
                   door interlock on power supply, we turn it on
                   at 400A.
            14:20 Try KAMI_KE3 configuration, rate of e's down
                   hole is very lower
            14:30 put ZCAL in beam, try KAMI_LAMPPI configuration,
                   rate of lambda's is even lower.
            15:00 zcal removed from beam, NM4AN off.
                   Rick Ford continues beam tuning.

 Priority for next shift continues to be beam tuning since we are
not yet sure how well beam is hitting the target.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:45:20 CST